EVA Report – February 5th

Crew 274 EVA Report 05-02-2023

EVA # 7

Author: Nicholas S. Pender, HSO

Purpose of EVA: EEVAC-1 (supply cache) data collection to ensure safe usage for future EVAs.

Start time: 09:30 AM

End time: 12:45 PM

Narrative: EVA #7’s purpose was to exercise an emergency scenario requiring the use of EEVAC-1 (the supply cache) located at coordinates N4248400, E519300. The MDRS hab location was used as a simulated location for an Earth supply drop that overshot it’s intended target, landing 3 hiking hours from base. In this scenario, two crew members (Nick and Bill) “drove” to the landing site (they didn’t actually drive to the landing site for this exercise), and upon arrival, their rover could not restart. Nick and Bill had to abandon the supply drop collection mission and return back to the hab on foot. At 9:35 AM, with an initial supply of 4 gel packs and 1.5L of water per person, Nick and Bill began their trek (from the MDRS) on foot, accompanied by the safety team (Sarah and Tony) on Perseverance. The safety team leap-frogged Nick and Bill ensuring positive line-of-sight and conducting occasional wellness checks. Consuming 1 gel pack every 30 minutes, averaging a pace of 2.4 mph, and with one rest stop along the way, Nick, Bill, and the safety team arrived at EEVAC-1 at 10:21 AM. Nick and Bill collected 3 additional gel packs and 1L of water from the cache while the safety team collected supply cache performance data. The water resupply required a maneuver that Nick and Bill rehearsed in space suits the previous evening, and it was executed exactly as rehearsed. At 10:32 AM, Nick and Bill resumed their trek South on Cow Dung Road, with the safety team following. Shortly after departing the cache, the charge on the safety team’s rover dropped to 61%. Since the remaining required distance beyond the cache was only 1 mile, the safety team decided to continue following Nick and Bill on foot, always keeping them in eyesight and comms range (the rover was parked around Zubrin’s Head at the bend and the Safety Team could not see the EVA team). It was during this period a break in sim occurred. A dog and his owner approached the safety team, and the dog became aggressive nipping Sarah on the leg. See HSO report for further details. With one final rest stop, the remaining scenario played out as intended, with Nick and Bill continuing the remaining 1.5 hours back to the MDRS hab. Their hike completed at 12:42 PM, and the crew reentered the airlock at 12:45.

*Note* There were multiple breaks in sim due to high vehicle traffic on Cow Dung Road.

Destination: South on Cow Dung Road.

Coordinates (use UTM WGS 84): N4247600, E520100

Participants: Nicholas Pender (HSO), Bill O’Hara (Journalist), Sarah Guthrie (Commander), and Tony DiBernardo (Journalist)

Road(s) and routes per MDRS Map: EVA #7’s (map provided) purpose is to drive to EEVAC-1 deployed at N4248400, E519300 via south on Cow Dung Road near Zubrin’s Head.

Mode of travel: Walking/Rover: Percy

Astronomy Report – February 5th

Astronomy Report

Name: Salina Peña and Noah Loy

Crew: 274

Date: 02/05/2023


Robotic Telescope Requested (Choose one) MDRS-14 MDRS-WF

Objects Viewed: None

Problems Encountered: No observations due to weather conditions (High winds and cloudy skies.)


Solar Features Observed: None

Problems Encountered: None due to cloud interference.

Operations Report – February 5th

Crew 274 Operations Report 2/5/2023

SOL: 7

Name of person filing report: Alexis Lojek

Non-nominal systems: Toilet.

Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews.


Spirit rover used: No

Hours: 208.0

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Opportunity rover used: No

Hours: 112.2

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Curiosity rover used: No

Hours: 220.3

Beginning charge: 100%

Ending charge: N/A

Currently charging: Yes

Perseverance rover used: Yes

Hours: 255.2

Beginning charge: 100%

Ending charge: 44%

Currently charging: No

General notes on rovers: Rover Perseverance was used during today’s EVA. The team returned after noon, so they were not plugged in.

Summary of Hab operations: Habitat is in a nominal condition.

WATER USE: 26 gallons as of 5:45 pm.

Water (static tank): 288 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged after EVA completion and taken off the chargers after the red charging lights turned off. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants from 2 – 2:40 pm.

WATER USE: 12.64 gallons used

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: No.

Summary of ScienceDome operations: GHO Officer Tyler Hines’ experiment is continuing to be monitored; Cress microgreens have made significant progress and are continuing to sprout.

Dual split: On a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: Based on our average water usage and the fact that we are a larger crew, we have calculated that the minimum water level of ~145 gallons (“31 inches” from the top of the lid) is likely to be reached on Sol 10. We would like a resupply of the 125 gallons of H2O, as offered at the beginning of our mission, as a safety net. Thank you!

GreenHab Report – February 5th

GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 80.75 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 12.68 gallons

Daily water usage for research and/or other purposes: 1.06 gallons

Water in Blue Tank: 100.05 gallons

Time(s) of watering for crops: 2:00 pm

Changes to crops: N/A

Narrative: Continued to conduct daily watering of plants and crops as part of daily required GreenHab operations. Following days of additional watering, tomato plants have shown a significant upturn in color and growth. In preparation for a potential final harvest and in parallel with daily operations, general cleaning and trimming of numerous plants and crops were also performed.

Harvest: N/A

Support/supplies needed: N/A

Journalist Report – February 5th

Crew 274 Journalist Report 2-5-2023

Sol: 7

Author: Tony DiBernardo, Crew Journalist

Title: The Long Hike

Crew 274 woke up this morning at 7:30am to the song “Take Me Home, Country Roads.” by John Denver. The crew got their surveys done, then made sure to get a decent breakfast before their 3-hour EVA.

EVA Activity

Today, HSO Pender and Hab Specialist O’Hara conducted a 3-hour EVA past the supply cache and back. The original EVA plan was to hike an hour to the supply cache, resupply their personal food and water from the cache, then hike 30 more minutes before turning around and hiking 1.5 hours back to the hab. They reached the cache early at the 47-minute mark due to them taking flat roads rather than the originally planned rocky terrain, but eventually found a more consistent and realistic pace on the way back. Supporting their EVA was Commander Guthrie and Journalist DiBernardo. They performed a “leap-frog” maneuver during the EVA, where they would drive the rover past the hiking crew and stop just before losing sight of them. Then, Pender and O’Hara would hike past the parked rover and just before they got out of site, Guthrie and DiBernardo would catch up in the rover and drive past them, only stopping just before losing sight again. This technique was used to reserve rover battery power as well as ensure a safe EVA for the crew on their longest hike of the mission. The EVA was a success, and apart from a couple of cars and hikers driving and passing by, the EVA went as expected.


Following the exciting germination of one LDEF seed yesterday, Greenhab Officer Hines, did his daily collection and observations of germination from the other seeds.. While no additional progress has been noted in other LDEF seeds, multiple sets of seeds that are yet to germinate were removed from the growth tent, placed in wet paper towels, and moved them into the GreenHab to utilize the higher temperatures and humidity in order to potentially germinate leading into the second half of the mission. Additionally, the cress microgreens have continued to make significant progress, with the red cabbage and broccoli making gradual yet minimal continued germination.


Crew astronomers Pena and Loy’s observations of heliophysics phenomena were significantly impaired by the site’s cloudy weather. Nevertheless, some useful data on solar dark spots and solar prominences were collected and analyzed. These observations will be very useful for the astronomer’s analysis of peak solar cycle 25. So far, 220,000 image points have been collected, and 50,000 images have been processed and analyzed.


Cloudy weather conditions canceled all planned observations of HADS Variable Star V0799 AUR. Crew Astronomers instead continued their analysis of their current image inventory, stacking and analyzing 12 separate 60-second images in the visual filter. So far, HADS Variable Star V0799 AUR is showing signs of dimming, which provides good data to the astronomers’ variable star report.

Stress Mitigation Study

Last night, the crew began their participation in Crew Engineer Lojeks’ study of tracking the stress levels of six of the crew members while practicing breathing and mindfulness techniques as a potential mitigation of stress in analog environments. He spoke with the crew about why he is pursuing this study, how to enter into this type of meditation and guided them through a 15-minute breathing exercise.

Sol Summary Report – February 5th

Crew 274 Sol Summary Report 02-03-2023

Sol: 7

Summary Title: Logistics Drop Successful

Author’s name: Sarah E. Guthrie (“Ceres”), Commander

Mission Status: Nominal

Sol Activity Summary: The crew completed EVA #7 which was a complex logistics cache scenario which involved the assessment of using the cache on a 3-hr hike for emergency rations. The test was a success and the crew was ecstatic with the results. While the research is in its first stages, the proof of concept leads to exciting larger applications for Lunar and Martian operations. The astronomy crew processed solar images earlier today, unfortunately the overcast that rolled in the evening has blocked night sky observations again. The crew closed the evening with debriefs, surveys, and prepped for the public outreach day scheduled for Monday (Sol 8).

Look Ahead Plan: The crew looks forward to its public outreach day tomorrow and speaking with elementary and middle school students, families, friends, and colleagues who have supported this mission and made it possible.

Anomalies in work: None

Weather: 3C High, -3C Low; Sunny until the evening, currently raining with 15mph winds

Crew Physical Status: Optimal

EVA: EVA #7 completed for logistics cache emergency scenario – successful

Reports to be filed: Sol Summary, Operations, HSO Report, EVA Report, GHO Report, and Astronomy Report.

Support Requested: Use of the MDRS Starlink to boost the signal for the public outreach event.

Copyright © The Mars Society. All rights reserved. | Main Site